# FairCLIP: Harnessing Fairness in Vision-and-Language Learning

### v1

@@ Abstract
Fairness is a critical concern in deep learning, especially in healthcare, where these models influence diagnoses and treatment decisions. Although fairness has been investigated in the vision-only domain, the fairness of medical vision-language (VL) models remains unexplored due to the scarcity of medical VL datasets for studying fairness. To bridge this research gap, we introduce the first fair vision-language medical dataset (FairVLMed) that provides detailed demographic attributes, ground-truth labels, and clinical notes to facilitate an in-depth examination of fairness within VL foundation models. Using FairVLMed, we conduct a comprehensive fairness analysis of two widely-used VL models (CLIP and BLIP2), pre-trained on both natural and medical domains, across four different protected attributes. Our results highlight significant biases in all VL models, with Asian, Male, Non-Hispanic, and Spanish being the preferred subgroups across the protected attributes of race, gender, ethnicity, and language, respectively. In order to alleviate these biases, we propose FairCLIP, an optimal-transport-based approach that achieves a favorable trade-off between performance and fairness by reducing the Sinkhorn distance between the overall sample distribution and the distributions corresponding to each demographic group. As the first VL dataset of its kind, FairVLMed holds the potential to catalyze advancements in the development of machine learning models that are both ethically aware and clinically effective. Our dataset and code are available at https://ophai.hms.harvard.edu/datasets/fairvlmed10k.
@@ 1 Introduction
Dataset Data Modality # of Images # of Patients Identity Attribute V L L Info Type Fitzpatrick17k [16] Skin Photos 16,012 1,373 Skin type ✓ ✗ HAM10000 [77] Dermatoscopy 9,948 - Age; Gender ✓ ✗ OL3I[92] Heart CT 8,139 1,686 Age; Gender ✓ ✗ ODIR-2019 [1] Color Fundus 8,000 5,000 Age; Gender ✓ ✗ PAPILA [35] Color Fundus 488 244 Age; Gender ✓ ✗ Harvard-GDP [47] OCT 1,000 1,000 Age; Gender; Race; Ethnicity ✓ ✗ COVID-CT-MD[3] Lung CT 308 305 Age; Gender ✓ ✗ AMD-OCT[14] OCT 384 - Age ✓ ✗ ADNI 1.5T [85] Brain MRI 550 Age; Gender ✓ ✗ CheXpert [24] Chest X-ray and Radiology Report 222,793 65,240 Age; Gender; Race ✓ ✓ Image Description MIMIC-CXR [27] Chest X-ray and Radiology Report 370,955 65,079 Age; Gender; Race ✓ ✓ Image Description PadChest [12] Chest X-ray and Radiology Report (Spanish) 160,868 69,882 Age; Gender ✓ ✓ Image Description FairVLMed SLO Fundus and Clinical Note Fundus: 10,000; Note: 10,000 10,000 Age; Gender; Race; Ethnicity; Preferred Language; Marital Status ✓ ✓ Image Description + Non-Imaging Clinical Info
@@
To bridge this research gap, we introduce the first fair vision-language medical dataset (FairVLMed for short) that provides detailed demographic attributes, ground-truth labels, and clinical notes to facilitate an in-depth examination of fairness within VL foundation models (Table 1). FairVLMed contains records for 10,000 patients, each paired with an SLO fundus image and a clinical note for diagnosing Glaucoma, along with fine-grained protected attributes such as age, gender, race, ethnicity, preferred language, and marital status. Unlike radiology reports, the clinical notes in our dataset provide much more detailed information featuring not only image descriptions but also rich non-imaging clinical information such as medication, non-imaging test results, and family history. Hence, these clinical notes are more representative of clinical textual information, making them better suited for studying the fairness of medical VL models.
Glaucoma, affecting millions globally [17, 74, 76, 49, 69, 72, 48, 71], exemplifies the need for equitable diagnostic models. Timely detection is crucial in order to avoid irreversible vision loss. However, many patients remain undiagnosed [68] due to the disease’s asymptomatic nature and barriers to ophthalmic care. Moreover, this undiagnosed issue is even more pronounced among minority groups. For example, previous studies have shown that individuals from Black communities are 4.4 more likely to have undiagnosed and untreated Glaucoma compared to their White counterparts [68], highlighting the importance of addressing healthcare disparities. Deep learning systems hold significant potential for improving healthcare. However, it is imperative to address potential fairness issues in these deep learning systems prior to their clinical implementation to ensure equitable healthcare delivery.
In this work, we conduct an extensive fairness analysis with two widely-used VL methods (i.e., CLIP [63] and BLIP2 [40]) on FairVLMed. Our experimental findings reveal significant disparities across various groups based on race, gender, ethnicity, and language. To address these fairness issues, we introduce FairCLIP, an optimal transport-based approach. FairCLIP is designed to enhance fairness by optimizing the Sinkhorn distance, thereby aligning the overall sample distribution with the distributions specific to each demographic group.
@@
We introduce the first fair vision-language medical dataset (FairVLMed) featuring detailed demographic attributes, ground-truth labels, and clinical notes for studying the fairness of VL foundation models.
Using FairVLMed, we conduct a comprehensive fairness analysis of two widely-used VL models (i.e., CLIP and BLIP2), pre-trained on both natural and medical domains, across four different protected attributes.
@@ 2 Related Work
Our FairVLMed dataset stands as a pioneering resource tailored for fairness studies in VL modeling. It comprises ground-truth labels, ensuring high label quality, and includes a broad range of identity attributes such as age, gender, race, ethnicity, preferred language, and marital status. Hence, this dataset not only enhances the quality and reliability of fairness studies but also broadens the horizon for research in the VL domain.
@@ 3.1 Data Collection and Quality Control
The subjects who received glaucoma services between 2015 and 2022 are from Massachusetts Eye and Ear at Harvard Medical School. There are three types of data to be released in this study: (1) scanning laser ophthalmoscopy (SLO) fundus images; (2) demographic identity group information; and (3) de-identified clinical notes written by ophthalmologists to provide a summary of the glaucoma diagnosis. SLO fundus images are a valuable marker for assessing retinal damage from diseases like glaucoma. Each SLO fundus image is associated with six demographic identity attributes, including age, gender, race, ethnicity, preferred language, and marital status. The accompanying clinical notes vary in length. These notes may detail assessments, treatment plans, and diagnostic strategies, and are considered to correspond with the visual semantics in SLO fundus images. Two examples of pairs of SLO fundus images and clinical notes are shown in Figure 1. The subjects are categorized into non-glaucoma (visual function measured by visual field (VF) test is normal: VF mean deviation -1 dB and normal VF glaucoma hemifield test and pattern standard deviation (PSD) results) and glaucoma categories (visual function measured by VF test is abnormal: VF mean deviation -3 dB and abnormal VF glaucoma hemifield test and PSD results).
@@ 3.3 Data Characteristics
The FairVLMed dataset comprises 10,000 samples from 10,000 subjects. It is divided into 7,000 training, 1,000 validation, and 2,000 test samples. The dataset’s collective age average is 60.9 16.2 years. The dataset includes samples from three major groups: Asian, with 819 samples; Black, with 1,491 samples; and White, with 7,690 samples. Gender-wise, females constitute 56.3% of the subjects, with the remainder being males. The ethnic distribution is highlighted by 90.6% Non-Hispanic, 4.0% Hispanic, and 5.4% unspecified. In terms of preferred language, 92.5% of the subjects prefer English, 1.7% prefer Spanish, 0.8% prefer other languages, and 5.0% remain unknown. From a marital status perspective, 57.4% are in a marriage or partnered, 26.4% are single, 6.6% have experienced divorce, 1.0% are legally separated, 6.1% are widowed, and 2.5% are not specified. After de-identification, clinical notes vary from 11 to 332 words, with an average word count of 147.
@@ 4.2 FairCLIP
where denotes the set of joint distributions whose first and second marginal distributions are and respectively, denotes the transport cost, and denotes the relative entropy of with respect to the product measure . and are the points in the distributions and , respectively. The Sinkhorn loss would be added to the loss used by CLIP in the pre-training phase for optimizing each component in CLIP.
@@ 5.1 Experimental Setup
In this section, we describe the pre-training and evaluation setups as well as the metrics used to study performance and fairness of VL models on our proposed FairVLMed dataset.
Pre-Training: We use the widely-adopted VL methods – CLIP [62] and BLIP2 [40] – for our analysis. For the natural pre-trained variants, we use the official checkpoints provided by CLIP and BLIP2. For the medical pre-trained variants, we pre-train both methods on the FairVLMed dataset after initializing from the official checkpoints. More details are in the supplementary material.
@@ 5.2 VL Fairness Analysis
Protected Attributes: Across the four protected attributes – race, gender, ethnicity, and language – our results indicate that VL models exhibit the best performance-fairness trade-off on ethnicity and the worst on language with average ES-AUC scores (across the four VL models) of 73.47 and 69.57 respectively. A granular analysis reveals that in terms of racial subgroups, Asian patients consistently have the highest diagnostic performance, whereas Black patients have the lowest. Across genders, male patients are consistently better diagnosed than female patients. Moreover, non-Hispanic is the highest-performing subgroup across ethnicities, whereas Spanish speakers have the best diagnostic performance across the language subgroups. Some of these performance disparities could be attributed to the imbalances in the pre-training datasets. For instance, non-Hispanic patients make up 90.6% of our dataset, potentially leading to improved performance in this subgroup. However, this is unlikely to be the only factor responsible for these performance disparities since the Asian, Male, and Spanish subgroups exhibit superior performances despite being the minority subgroups. This indicates that the pre-training of these models could potentially play a role in the biases exhibited by these models. We delve deeper into this in the subsequent analysis.
@@
Natural vs. Medical Pre-training: Here, we compare the VL models pre-trained on two different domains – natural vs. medical VL datasets. Comparing the two pre-training domains, we observe that pre-training on medical data (i.e., paired fundus images + clinical notes) improves the performance-fairness trade-off across all protected attributes except language. Concretely, we see the most benefit on ethnicity, with improvements of as much as 8.38 in terms of ES-AUC. A granular analysis of the individual subgroups shows that the White and Hispanic subgroups witness the most improvement from medical pre-training, with improvements of as much as 6.75 and 9.65 ES-AUC in the former and latter respectively. This indicates that medical VL pre-training could be an effective strategy for improving the performances of under-performing subgroups, thereby effectively reducing the performance disparities across different protected attributes.
@@ 5.4 Ablation Studies
Clinical Note Summarization: VL models are usually trained with images and captions, which tend to be quite short. In contrast, the clinical notes in our FairVLMed dataset are fairly lengthy, capturing a lot more nuanced information. Hence, we summarize these clinical notes in order to align our setup with the standard VL frameworks while retaining essential medical information. We use three different types of LLMs for summarization. In addition to the SOTA GPT-4 [56] model, we also use two LLMs specialized for the medical domain – PMC-LLAMA [82] and MED42222https://huggingface.co/m42-health/med42-70b. We use the following prompt for summarization for all three LLMs: Summarize the key details, including the presence of glaucoma, from the clinical note within 180 characters. Figure 3(a) compares the performance-fairness trade-off of BLIP2 models trained via the three summarized notes against the original notes. Across all protected attributes except language, we note that all three summarization methods yield a consistent improvement in ES-AUC. Particularly, MED42 yields the best performance across race, GPT-4 yields the best performance across gender, whereas PMC-LLAMA yields the best performance across both ethnicity and language. Overall, the medical pre-trained LLMs (PMC-LLAMA and MED42) yield superior results, highlighting the efficacy of domain-specific LLMs for clinical note summarization.
@@ 6 Conclusion
Given the critical need for fairness in healthcare-focused settings, we introduce the first fair vision-language medical dataset (FairVLMed) for studying the fairness of medical VL foundation models. Our comprehensive fairness analysis on FairVLMed reveals significant biases in all VL models. To address these biases, we propose FairCLIP, an optimal transport-based approach that effectively balances both performance and fairness. To aid future research in this area, we make our dataset and code publicly available at https://ophai.hms.harvard.edu/datasets/fairvlmed10k.

### v2

# FairCLIP: Harnessing Fairness in Vision-Language Learning

@@ Abstract
Fairness is a critical concern in deep learning, especially in healthcare, where these models influence diagnoses and treatment decisions. Although fairness has been investigated in the vision-only domain, the fairness of medical vision-language (VL) models remains unexplored due to the scarcity of medical VL datasets for studying fairness. To bridge this research gap, we introduce the first fair vision-language medical dataset (Harvard-FairVLMed) that provides detailed demographic attributes, ground-truth labels, and clinical notes to facilitate an in-depth examination of fairness within VL foundation models. Using Harvard-FairVLMed, we conduct a comprehensive fairness analysis of two widely-used VL models (CLIP and BLIP2), pre-trained on both natural and medical domains, across four different protected attributes. Our results highlight significant biases in all VL models, with Asian, Male, Non-Hispanic, and Spanish being the preferred subgroups across the protected attributes of race, gender, ethnicity, and language, respectively. In order to alleviate these biases, we propose FairCLIP, an optimal-transport-based approach that achieves a favorable trade-off between performance and fairness by reducing the Sinkhorn distance between the overall sample distribution and the distributions corresponding to each demographic group. As the first VL dataset of its kind, Harvard-FairVLMed holds the potential to catalyze advancements in the development of machine learning models that are both ethically aware and clinically effective. Our dataset and code are available at https://ophai.hms.harvard.edu/datasets/harvard-fairvlmed10k.
@@ 1 Introduction
Dataset Data Modality # of Images # of Patients Identity Attribute V L L Info Type Fitzpatrick17k [16] Skin Photos 16,012 1,373 Skin type ✓ ✗ HAM10000 [77] Dermatoscopy 9,948 - Age; Gender ✓ ✗ OL3I[92] Heart CT 8,139 1,686 Age; Gender ✓ ✗ ODIR-2019 [1] Color Fundus 8,000 5,000 Age; Gender ✓ ✗ PAPILA [35] Color Fundus 488 244 Age; Gender ✓ ✗ Harvard-GDP [47] OCT 1,000 1,000 Age; Gender; Race; Ethnicity ✓ ✗ COVID-CT-MD[3] Lung CT 308 305 Age; Gender ✓ ✗ AMD-OCT[14] OCT 384 - Age ✓ ✗ ADNI 1.5T [85] Brain MRI 550 Age; Gender ✓ ✗ CheXpert [24] Chest X-ray and Radiology Report 222,793 65,240 Age; Gender; Race ✓ ✓ Image Description MIMIC-CXR [27] Chest X-ray and Radiology Report 370,955 65,079 Age; Gender; Race ✓ ✓ Image Description PadChest [12] Chest X-ray and Radiology Report (Spanish) 160,868 69,882 Age; Gender ✓ ✓ Image Description Harvard-FairVLMed SLO Fundus and Clinical Note Fundus: 10,000; Note: 10,000 10,000 Age; Gender; Race; Ethnicity; Preferred Language; Marital Status ✓ ✓ Image Description + Non-Imaging Clinical Info
@@
To bridge this research gap, we introduce the first fair vision-language medical dataset (Harvard-FairVLMed for short) that provides detailed demographic attributes, ground-truth labels, and clinical notes to facilitate an in-depth examination of fairness within VL foundation models (Table 1). Harvard-FairVLMed contains records for 10,000 patients, each paired with an SLO fundus image and a clinical note for diagnosing Glaucoma, along with fine-grained protected attributes such as age, gender, race, ethnicity, preferred language, and marital status. Unlike radiology reports, the clinical notes in our dataset provide much more detailed information featuring not only image descriptions but also rich non-imaging clinical information such as medication, non-imaging test results, and family history. Hence, these clinical notes are more representative of clinical textual information, making them better suited for studying the fairness of medical VL models.
Glaucoma, affecting millions globally [17, 74, 76, 49, 69, 72, 48, 71], exemplifies the need for equitable diagnostic models. Timely detection is crucial in order to avoid irreversible vision loss. However, many patients remain undiagnosed [68] due to the disease’s asymptomatic nature and barriers to ophthalmic care. Moreover, this undiagnosed issue is even more pronounced among minority groups. For example, previous studies have shown that individuals from Black communities are 4.4 times more likely to have undiagnosed and untreated Glaucoma compared to their White counterparts [68], highlighting the importance of addressing healthcare disparities. Deep learning systems hold significant potential for improving healthcare. However, it is imperative to address potential fairness issues in these deep learning systems prior to their clinical implementation to ensure equitable healthcare delivery.
In this work, we conduct an extensive fairness analysis with two widely-used VL methods (i.e., CLIP [63] and BLIP2 [40]) on Harvard-FairVLMed. Our experimental findings reveal significant disparities across various groups based on race, gender, ethnicity, and language. To address these fairness issues, we introduce FairCLIP, an optimal transport-based approach. FairCLIP is designed to enhance fairness by optimizing the Sinkhorn distance, thereby aligning the overall sample distribution with the distributions specific to each demographic group.
@@
We introduce the first fair vision-language medical dataset (Harvard-FairVLMed) featuring detailed demographic attributes, ground-truth labels, and clinical notes for studying the fairness of VL foundation models.
Using Harvard-FairVLMed, we conduct a comprehensive fairness analysis of two widely-used VL models (i.e., CLIP and BLIP2), pre-trained on both natural and medical domains, across four different protected attributes.
@@ 2 Related Work
Our Harvard-FairVLMed dataset stands as a pioneering resource tailored for fairness studies in VL modeling. It comprises ground-truth labels, ensuring high label quality, and includes a broad range of identity attributes such as age, gender, race, ethnicity, preferred language, and marital status. Hence, this dataset not only enhances the quality and reliability of fairness studies but also broadens the horizon for research in the VL domain.
@@ 3.1 Data Collection and Quality Control
Harvard-FairVLMed contains subjects who received glaucoma care from Massachusetts Eye and Ear at Harvard Medical School between 2015 and 2022. There are three types of data to be released in this study: (1) scanning laser ophthalmoscopy (SLO) fundus images; (2) demographic identity group information; and (3) de-identified clinical notes written by ophthalmologists to provide a summary of the glaucoma diagnosis. SLO fundus images are a valuable marker for assessing retinal damage from diseases like glaucoma. Each SLO fundus image is associated with six demographic identity attributes, including age, gender, race, ethnicity, preferred language, and marital status. The accompanying clinical notes vary in length. These notes may detail assessments, treatment plans, and diagnostic strategies, and are considered to correspond with the visual semantics in SLO fundus images. Two examples of pairs of SLO fundus images and clinical notes are shown in Figure 1. The subjects are categorized into non-glaucoma (visual function measured by visual field (VF) test is normal: VF mean deviation -1 dB and normal VF glaucoma hemifield test and pattern standard deviation (PSD) results) and glaucoma categories (visual function measured by VF test is abnormal: VF mean deviation -3 dB and abnormal VF glaucoma hemifield test and PSD results).
@@ 3.3 Data Characteristics
Our Harvard-FairVLMed dataset comprises 10,000 samples from 10,000 subjects. It is divided into 7,000 training, 1,000 validation, and 2,000 test samples. The dataset’s collective age average is 60.9 16.2 years. The dataset includes samples from three major groups: Asian, with 819 samples; Black, with 1,491 samples; and White, with 7,690 samples. Gender-wise, females constitute 56.3% of the subjects, with the remainder being males. The ethnic distribution is highlighted by 90.6% Non-Hispanic, 4.0% Hispanic, and 5.4% unspecified. In terms of preferred language, 92.5% of the subjects prefer English, 1.7% prefer Spanish, 0.8% prefer other languages, and 5.0% remain unknown. From a marital status perspective, 57.4% are in a marriage or partnered, 26.4% are single, 6.6% have experienced divorce, 1.0% are legally separated, 6.1% are widowed, and 2.5% are not specified. After de-identification, clinical notes vary from 11 to 332 words, with an average word count of 147.
@@ 4.2 FairCLIP
where denotes the set of joint distributions whose first and second marginal distributions are and , respectively, denotes the transport cost, and denotes the relative entropy of with respect to the product measure . and are the points in the distributions and , respectively. The Sinkhorn loss would be added to the loss used by CLIP in the pre-training phase for optimizing each component in CLIP.
@@ 5.1 Experimental Setup
In this section, we describe the pre-training and evaluation setups as well as the metrics used to study performance and fairness of VL models on our proposed Harvard-FairVLMed dataset.
Pre-Training: We use the widely-adopted VL methods – CLIP [62] and BLIP2 [40] – for our analysis. For the natural pre-trained variants, we use the official checkpoints provided by CLIP and BLIP2. For the medical pre-trained variants, we pre-train both methods on the Harvard-FairVLMed dataset after initializing from the official checkpoints. More details are in the supplementary material.
@@ 5.2 VL Fairness Analysis
Protected Attributes: Across the four protected attributes – race, gender, ethnicity, and language – our results indicate that VL models exhibit the best performance-fairness trade-off on ethnicity and the worst on language with average ES-AUC scores (across the four VL models) of 73.47 and 69.57, respectively. A granular analysis reveals that in terms of racial subgroups, Asian patients consistently have the highest diagnostic performance, whereas Black patients have the lowest. Across genders, male patients are consistently better diagnosed than female patients. Moreover, non-Hispanic is the highest-performing subgroup across ethnicities, whereas Spanish speakers have the best diagnostic performance across the language subgroups. Some of these performance disparities could be attributed to the imbalances in the pre-training datasets. For instance, non-Hispanic patients make up 90.6% of our dataset, potentially leading to improved performance in this subgroup. However, this is unlikely to be the only factor responsible for these performance disparities since the Asian, Male, and Spanish subgroups exhibit superior performances despite being the minority subgroups. This indicates that the pre-training of these models could potentially play a role in the biases exhibited by these models. We delve deeper into this in the subsequent analysis.
@@
Natural vs. Medical Pre-training: Here, we compare the VL models pre-trained on two different domains – natural vs. medical VL datasets. Comparing the two pre-training domains, we observe that pre-training on medical data (i.e., paired fundus images + clinical notes) improves the performance-fairness trade-off across all protected attributes except language. Concretely, we see the most benefit on ethnicity, with improvements of as much as 8.38 in terms of ES-AUC. A granular analysis of the individual subgroups shows that the White and Hispanic subgroups witness the most improvement from medical pre-training, with improvements of as much as 6.75 and 9.65 ES-AUC in the former and latter, respectively. This indicates that medical VL pre-training could be an effective strategy for improving the performances of under-performing subgroups, thereby effectively reducing the performance disparities across different protected attributes.
@@ 5.4 Ablation Studies
Clinical Note Summarization: VL models are usually trained with images and captions, which tend to be quite short. In contrast, the clinical notes in our Harvard-FairVLMed dataset are fairly lengthy, capturing a lot more nuanced information. Hence, we summarize these clinical notes in order to align our setup with the standard VL frameworks while retaining essential medical information. We use three different types of LLMs for summarization. In addition to the SOTA GPT-4 [56] model, we also use two LLMs specialized for the medical domain – PMC-LLAMA [82] and MED42222https://huggingface.co/m42-health/med42-70b. We use the following prompt for summarization for all three LLMs: Summarize the key details, including the presence of glaucoma, from the clinical note within 180 characters. Figure 2(a) compares the performance-fairness trade-off of BLIP2 models trained via the three summarized notes against the original notes. Across all protected attributes except language, we note that all three summarization methods yield a consistent improvement in ES-AUC. Particularly, MED42 yields the best performance across race, GPT-4 yields the best performance across gender, whereas PMC-LLAMA yields the best performance across both ethnicity and language. Overall, the medical pre-trained LLMs (PMC-LLAMA and MED42) yield superior results, highlighting the efficacy of domain-specific LLMs for clinical note summarization.
@@ 6 Conclusion
Given the critical need for fairness in healthcare-focused settings, we introduce the first fair vision-language medical dataset (Harvard-FairVLMed) for studying the fairness of medical VL foundation models. Our comprehensive fairness analysis on Harvard-FairVLMed reveals significant biases in all VL models. To address these biases, we propose FairCLIP, an optimal transport-based approach that effectively balances both performance and fairness. To aid future research in this area, we make our dataset and code publicly available at https://ophai.hms.harvard.edu/datasets/harvard-fairvlmed10k.